# Enhancing Step-by-Step and Verifiable Medical Reasoning in MLLMs

### v1

@@ Abstract
Multimodal large language models (MLLMs) have begun to demonstrate robust reasoning capabilities on general tasks, yet their application in the medical domain remains in its early stages. Constructing chain-of-thought (CoT) training data is essential for bolstering the reasoning abilities of medical MLLMs. However, existing approaches exhibit a deficiency in offering a comprehensive framework for searching and evaluating effective reasoning paths towards critical diagnosis. To address this challenge, we propose Mentor-Intern Collaborative Search (MICS), a novel reasoning-path searching scheme to generate rigorous and effective medical CoT data. MICS first leverages mentor models to initialize the reasoning, one step at a time, then prompts each intern model to continue the thinking along those initiated paths, and finally selects the optimal reasoning path according to the overall reasoning performance of multiple intern models. The reasoning performance is determined by an MICS-Score, which assesses the quality of generated reasoning paths. Eventually, we construct MMRP, a multi-task medical reasoning dataset with ranked difficulty, and Chiron-o1, a new medical MLLM devised via a curriculum learning strategy, with robust visual question-answering and generalizable reasoning capabilities. Extensive experiments demonstrate that Chiron-o1, trained on our CoT dataset constructed using MICS, achieves state-of-the-art performance across a list of medical visual question answering and reasoning benchmarks. Code will be available at https://github.com/manglu097/Chiron-o1
11footnotetext: Equal contribution. 🖂 Corresponding authors. This Work is under support from the Shanghai Artificial Intelligence Laboratory.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) chen2024far ; hurst2024gpt ; laurenccon2024building ; liu2023visual ; team2023gemini ; yao2024minicpm have demonstrated prominent performance in a wide range of tasks, including image captioning, visual question answering, and video analysis. Recent breakthroughs in MLLMs have also shed new light on the development of general-purpose medical AI. Remarkable efforts have been made to adapt existing MLLMs to address complex clinical tasks through supervised fine-tuning (SFT) on carefully curated multimodal medical instruction fine-tuning datasets chen2024huatuogpt ; li2023llava ; li2024gmai ; wu2024pmc ; yang2024advancing . However, most current medical MLLMs rely on a direct-prediction paradigm, which produces brief, immediate answers to problems and often overlooks the interleaved image–text reasoning processes essential for real-world clinical scenarios. This critical shortcoming—the inability to perform deep, multimodal reasoning analysis of medical data—has given rise to an urgent need for the development of more sophisticated medical MLLMs capable of tackling complex clinical challenges and offering improved diagnostic support.
Building on recent advancements in reinforcement learning (RL) optimization, several promising approaches lai2025med ; pan2025medvlm ; su2025gmai have leveraged Group Relative Policy Optimization (GRPO) shao2024deepseekmath to elicit reasoning capabilities in models, introducing multiple R1-series medical MLLMs. Nevertheless, these models continue to exhibit limited performance in answering real-world clinical questions. As pointed out by recent studies yue2025does , although RL training improves performance by biasing the model’s output distribution toward reward-yielding trajectories, it fails to generate novel reasoning paradigms. Consequently, the performance ceiling of the R1 models has long been constrained by their underlying base MLLM. One potential solution is to incorporate high-quality chain-of-thought (CoT) annotations into SFT to bolster the model’s reasoning capabilities and foster novel reasoning paradigms. However, unlike general-domain tasks in which large-scale CoT datasets can be crowdsourced, medical reasoning demands domain-specific logical structures and clinical expertise. Curating such medical CoT datasets is prohibitively expensive and time-consuming. Furthermore, the absence of standardized evaluation metrics to ensure the validity of generated CoT processes also remains a critical barrier to step-by-step visual reasoning annotation. Therefore, it is crucial to develop an effective and sophisticated method that can generate intermediate reasoning steps toward the final answer and evaluate the step-by-step visual reasoning quality.
@@
2 Related Works
Reasoning in Medical MLLMs. As multimodal reasoning demand grows, frameworks and techniques have evolved, extending the reasoning capabilities of Large Language Models (LLMs) guo2025deepseek ; team2025kimi ; qwq32b to tackle more complex general-purpose vision-language tasks touvron2023llama ; yao2024minicpm ; chen2024internvl ; bai2025qwen2 . Techniques such as prompt tuning zamfirescu2023johnny , SFT shen2024rethinking , and RL chu2025sft have emerged as critical methods for boosting the reasoning capabilities of MLLMs. Given its potential to enhance model’s performance on complex tasks, reasoning ability has also garnered significant attention in medical domains chen2024huatuogpt ; wang2025baichuan . However, equipping MLLMs with reasoning capabilities by integrating multimodal medical information remains an open and challenging problem. Recent studies liu2024medcot ; wei2024mc have explored designing well-crafted prompts to simulate the reasoning process from medical problems to correct answers, often involving multiple doctor roles or functional modules. However, the applicability of the aforementioned methods is limited. Meanwhile, as DeepSeek-R1 significantly enhances model reasoning capabilities through GRPO guo2025deepseek , an increasing number of studies follow this paradigm to adapt RL to the medical domain. Med-R1 lai2025med , utilizing the GRPO strategy, demonstrates superior performance across different image modalities, enhancing the generalizability of MLLMs in medical reasoning. Similarly, MedVLM-R1 pan2025medvlm adopts an RL framework to encourage the discovery of human-interpretable reasoning paths. However, these RL-based models rely solely on fixed reward functions (e.g., format and accuracy), neglecting the evaluation of reasoning paths, which can easily lead to superficial or hallucinated reasoning processes. In contrast, our proposed method leverages a collaborative search to identify high-quality reasoning paths with minimal hallucinations, thereby enhancing reasoning abilities.
@@ 2 Related Works
Construction of CoT Datasets. The development of CoT datasets for SFT is crucial in multimodal reasoning, setting it apart from traditional models alayrac2022flamingo ; li2023blip ; zhu2023minigpt ; liu2024improved by emphasizing the reasoning process rather than merely providing final answers wang2022self ; wei2022chain . Research in this field can be categorized into prompt-based, plan-based, and learning-based methods wang2025multimodal , each contributing uniquely to the CoT generation in multimodal contexts. Prompt-based methods utilize carefully designed prompts to guide models in generating CoT during inference. Simple prompts, such as "think step-by-step" himakunthala2023let , initiate creation of reasoning paths for multimodal tasks, while advanced approaches establish a clear reasoning workflow by specifying detailed objectives and procedures for each step chen2023see ; fei2024video ; meng2023chain . Plan-based methods enable dynamic exploration of reasoning paths, enhancing adaptability. For example, MM-ToT kye2023mmtot combines GPT-4 hurst2024gpt and Stable Diffusion rombach2021highresolution , using depth-first and breadth-first searches to select optimal solutions. Additionally, learning-based methods embed CoT construction within model training or fine-tuning. PCoT wang2024t optimizes this approach for CoT generation, whereas MC-CoT tan2024boosting improves reasoning capabilities in smaller-scale models via majority voting during training. However, as constructing medical CoT requires specific logic and domain expertise, the aforementioned methods cannot be effectively transferred to the medical field. In other words, the construction of multimodal CoT in the medical domain remains an underexplored stage. Existing studies wei2024mc primarily rely on manual annotations or model-generated reasoning paths without any evaluation. Our proposed method takes a step further by addressing these challenges by leveraging an automatic and collaborative search strategy to identify high-quality reasoning paths.
@@ 3.2 Mentor-Intern Collaborative Search for Effective Reasoning
where $\text{MICS}_{k}$ denotes the search process up to the step $k$, and $\{s_{1,x},s_{2,y},...,s_{k,z}\}$ represents the all reasoning steps searched thus far, serving as the reasoning path prefix for the subsequent search. $P$, $Q$, $A$, and $I$ represent patient information, questions, ground truth, and corresponding images.
@@
[…]
@@ 4.1 Experiment Settings
In our experiments, the MICS strategy employs three mentor models: ChatGPT-4o hurst2024gpt , Gemini 2.5 Pro Preview team2024gemini , and Qwen2.5-VL-72B-Instruct bai2025qwen2 . After careful selection, we designate three open-source models as intern models, each with two distinct temperature (0.3 and 1.2): Qwen2-VL-7B wang2024qwen2 , Qwen2.5-VL-7B bai2025qwen2 , and InternVL3-8B chen2024internvl . We utilize DeepSeek-V3 liu2024deepseek as the "judge" to evaluate the alignment between the answers generated by intern models and the ground truth. To balance search accuracy and efficiency, the maximum search depth is set to 4. During the training phase, we adopt InternVL3-8B as the base model for Chiron-o1, employing LoRA for fine-tuning hu2022lora , with AdamW as the optimizer, a learning rate of 4e-5 following a cosine decay schedule, a batch size of 16, and bfloat16 mixed precision. The training data comprise the MMRP dataset and commonly-used VQA datasets, with details provided in the Appendix C.
@@
We compare Chiron-o1 with the following three types of baseline models: (1) General MLLMs, including high-performance general vision models, both closed-source (GPT series hurst2024gpt , Gemini series team2024gemini ) and open-source (LLaVA series liu2024improved , Qwen series bai2025qwen2 ). (2) Medical MLLMs, comprising vision models pretrained on specific medical corpora, such as LLaVA-Med li2023llava , Med-Flamingo moor2023med , GMAI-VL li2024gmai , and HuatuoGPT-Vision chen2024huatuogpt . (3) Medical Reasoning MLLMs, including Med-R1 lai2025med and MedVLM-R1pan2025medvlm , which are fine-tuned using GRPO shao2024deepseekmath . For evaluation metrics, we use choice accuracy to measure model performance on VQA benchmarks. For the open-ended questions requiring reasoning in Part 3 of the MMRP dataset, BERT-Score zhang2019bertscore is employed to assess the semantic similarity between the model’s final answer and the ground truth, while MICS-Score is used to evaluate the effectiveness of the reasoning path.
4.2 Main Results
Performance on the Medical VQA Benchmarks
Medical models should strive to further enhance their performance on VQA tasks while improving reasoning capabilities. Therefore, to evaluate the foundational performance of Chiron-o1, we first test it on several commonly-used VQA benchmarks. The results, compared against other mainstream models, are summarized in Table 1. Compared to the baseline models, Chiron-o1 achieves significant performance improvements across five benchmarks. Specifically, Chiron-o1-8B and Chiron-o1-2B outperform their respective baseline models by an average of 5.7% and 8.1%. This demonstrates that our reasoning-focused model further enhances visual understanding and question-answering capabilities. Subsequently, Chiron-o1 outperforms most medical MLLMs across all benchmarks. It even achieves comparable performance to HuatuoGPT-Vision-34B, which has four times the parameters, on the PMC-VQA and MMMU (Health & Medicine) benchmarks, while significantly surpassing it on the remaining benchmarks, with an overall average improvement of 5%. Finally, we observe that existing medical reasoning models, such as Med-R1 and MedVLM-R1, tend to lose basic VQA capabilities while focusing on reasoning, as shown in Table 1. In contrast, Chiron-o1-8B outperforms them on VQA benchmarks by an average of 20.6% and 16.6%, respectively. Overall, Chiron-o1 demonstrates competitive performance across multiple benchmarks, highlighting its versatility in medical image understanding and question answering.
@@ 4.2 Main Results
Gemini Pro team2023gemini 60.3 72.6 70.3 47.9 62.8
@@ 4.3 Ablation Studies
We apply the MICS strategy to build the MMRP multi-task dataset to improve the model’s multimodal reasoning. An ablation study on Chiron-o1-8B evaluates its impact. Unlike MICS, the vanilla method enables the mentor model to generate reasoning directly, bypassing the evaluation of intern models. Table 3 shows that excluding reasoning data (first row) or using reasoning paths constructed by vanilla method(second and third rows) degrades performance. Compared to Chiron-o1-8B, the former reduces average performance by 3% across four benchmarks, and the latter by 2.7%. Furthermore, Figure 4(a) illustrates the differences in effectiveness between reasoning paths constructed using MICS and the vanilla method across three distinct medical scenarios. We compute the MICS-Score step-by-step for an entire reasoning path to evaluate the value of reasoning prefixes and analyze the trend of MICS-Score changes (categorized into four types, with a monotonically increasing score indicating an effective trend and the others deemed low-value, details provided in the Appendix B). As depicted in Figure 4(a), the proportion of effective reasoning paths identified by MICS significantly surpasses that of the vanilla method. These results underscore the critical role of the MICS strategy in searching for effective reasoning paths.
@@
Figure 4: Ablation Studies on MICS and Model Training. (a) Contribution of the MICS strategy to reasoning path score trends, with a, b, and c denoting three clinical scenarios (Appendix A.3). "vanilla" refers to directly generating reasoning paths using the mentor model without evaluation. (b) Investigation of the advantages of our training strategy compared to other training methods.
@@
During training, we add VQA datasets to enhance the model’s visual understanding. To assess their impact, we compare a model trained without VQA datasets to Chiron-o1-8B. Table 3 (second and fourth rows) shows VQA data improves performance on simpler benchmarks, with VQA-RAD and SLAKE increasing by 5% and 7.3%, respectively, but has limited effect on complex reasoning benchmarks. We also examine whether quality control of MMRP reasoning data boosts reasoning performance. Table 3 (fifth row) indicates quality control raises accuracy on MMMU (H&M) from 49.3% to 54.6% (+5.3%) and on MedXpertQA_MM from 22.6% to 24.2% (+1.6%). These results confirm the effectiveness of our train set configuration.
@@
In Section 3.3, we introduce a stage-wise training strategy for model fine-tuning. To assess its impact, we traine models using stage subsets and compare their performance against Chiron-o1-8B on multiple benchmarks. For the complex VQA benchmark, such as MMMU (H&M), Figure 4(b) shows that models trained with only Stage 1, Stage 1+2, or Stage 3 exhibit performance drops of 6.9%, 6.2%, and 3.5%, respectively, compared to Chiron-o1-8B. Omitting Stage 3 notably weakens reasoning capabilities, with models trained on Stage 1 or Stage 1+2 scoring 3.1% and 0.25% lower on average than Stage 3 alone on MMMU (H&M) and MedXpertQA_MM. Similarly, training only with Stage 3, without enhancing visual question-answering, significantly reduces performance on VQA-RAD and SLAKE compared to other ablation models.

### v2

# Chiron-o1: Igniting Multimodal Large Language Models towards Generalizable Medical Reasoning via Mentor-Intern Collaborative Search

@@ Abstract
Multimodal large language models (MLLMs) have begun to demonstrate robust reasoning capabilities on general tasks, yet their application in the medical domain remains in its early stages. Constructing chain-of-thought (CoT) training data is essential for bolstering the reasoning abilities of medical MLLMs. However, existing approaches exhibit a deficiency in offering a comprehensive framework for searching and evaluating effective reasoning paths towards critical diagnosis. To address this challenge, we propose Mentor-Intern Collaborative Search (MICS), a novel reasoning-path searching scheme to generate rigorous and effective medical CoT data. MICS first leverages mentor models to initialize the reasoning, one step at a time, then prompts each intern model to continue the thinking along those initiated paths, and finally selects the optimal reasoning path according to the overall reasoning performance of multiple intern models. The reasoning performance is determined by an MICS-Score, which assesses the quality of generated reasoning paths. Eventually, we construct MMRP, a multi-task medical reasoning dataset with ranked difficulty, and Chiron-o1, a new medical MLLM devised via a curriculum learning strategy, with robust visual question-answering and generalizable reasoning capabilities. Extensive experiments demonstrate that Chiron-o1, trained on our CoT dataset constructed using MICS, achieves state-of-the-art performance across a list of medical visual question answering and reasoning benchmarks. Codes are available at https://github.com/manglu097/Chiron-o1
11footnotetext: Equal contribution. 🖂 Corresponding authors.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) chen2024far ; hurst2024gpt ; laurenccon2024building ; liu2023visual ; team2023gemini ; yao2024minicpm have demonstrated prominent performance in a wide range of tasks, including image captioning, visual question answering, and video analysis. Recent breakthroughs in MLLMs have also shed new light on the development of general‑purpose medical AI. Remarkable efforts have been made to adapt existing MLLMs to address complex clinical tasks through supervised fine‑tuning (SFT) on carefully curated multimodal medical instruction fine-tuning datasets chen2024huatuogpt ; li2023llava ; li2024gmai ; wu2024pmc ; yang2024advancing . However, most current medical MLLMs rely on a direct‑prediction paradigm, which produces brief, immediate answers to problems and often overlooks the interleaved image–text reasoning processes essential for real‑world clinical scenarios. This critical shortcoming, namely the inability to perform deep multimodal reasoning analysis of medical data, has given rise to an urgent need for the development of more sophisticated medical MLLMs capable of tackling complex clinical challenges and offering improved diagnostic support.
Building on recent advancements in reinforcement learning (RL) optimization, several promising approaches lai2025med ; pan2025medvlm ; su2025gmai have leveraged Group Relative Policy Optimization (GRPO) shao2024deepseekmath to elicit reasoning capabilities in models, introducing multiple R1‑series medical MLLMs. Nevertheless, these models continue to exhibit limited performance in answering real-world clinical questions. As pointed out by recent studies yue2025does , although RL training improves performance by biasing the model’s output distribution toward reward‑yielding trajectories, it fails to generate novel reasoning paradigms. Consequently, the performance ceiling of the R1 models has long been constrained by their underlying base MLLM. One potential solution is to incorporate high-quality chain-of-thought (CoT) annotations into SFT to bolster the model’s reasoning capabilities and foster novel reasoning paradigms. However, unlike general‑domain tasks in which large‑scale CoT datasets can be crowdsourced, medical reasoning demands domain‑specific logical structures and clinical expertise. Curating such medical CoT datasets is prohibitively expensive and time‑consuming. Furthermore, the absence of standardized evaluation metrics to ensure the validity of generated CoT processes also remains a critical barrier to step-by-step visual reasoning annotation. Therefore, it is crucial to develop an effective and sophisticated method that can generate intermediate reasoning steps toward the final answer and evaluate the step-by-step visual reasoning quality.
@@
2 Related Works
Reasoning in Medical MLLMs. As multimodal reasoning demand grows, frameworks and techniques have evolved, extending the reasoning capabilities of Large Language Models (LLMs) guo2025deepseek ; team2025kimi ; qwq32b to tackle more complex general-purpose vision-language tasks touvron2023llama ; yao2024minicpm ; chen2024internvl ; bai2025qwen2 . Techniques such as prompt tuning zamfirescu2023johnny , SFT shen2024rethinking , and RL chu2025sft have emerged as critical methods for boosting the reasoning capabilities of MLLMs. Given its potential to enhance model’s performance on complex tasks, reasoning ability has also garnered significant attention in medical domains. Models like HuatuoGPT-o1 chen2024huatuogpt and Baichuan-M1 wang2025baichuan are dedicated to enhancing the ability to solve medical problems through rigorous and transparent reasoning paths. However, equipping MLLMs with reasoning capabilities by integrating multimodal medical information remains an open and challenging problem. Recent studies liu2024medcot ; wei2024mc have explored designing well-crafted prompts to simulate the reasoning process from medical problems to correct answers, often involving multiple doctor roles or functional modules. However, the applicability of the aforementioned methods is limited. Meanwhile, as DeepSeek-R1 significantly enhances model reasoning capabilities through GRPO guo2025deepseek , an increasing number of studies follow this paradigm to adapt RL to the medical domain. Med-R1 lai2025med , utilizing the GRPO strategy, demonstrates superior performance across different image modalities, enhancing the generalizability of MLLMs in medical reasoning. Similarly, MedVLM-R1 pan2025medvlm adopts an RL framework to encourage the discovery of human-interpretable reasoning paths. However, these RL-based models rely solely on fixed reward functions (e.g., format and accuracy), neglecting the evaluation of reasoning paths, which can easily lead to superficial or hallucinated reasoning processes. In contrast, our proposed method leverages a collaborative search to identify high-quality reasoning paths with minimal hallucinations, thereby enhancing reasoning abilities.
@@ 2 Related Works
Construction of CoT Datasets. The development of CoT datasets for SFT is crucial in multimodal reasoning, setting it apart from traditional models alayrac2022flamingo ; li2023blip ; zhu2023minigpt ; liu2024improved by emphasizing the reasoning process rather than merely providing final answers wang2022self ; wei2022chain . Research in this field can be categorized into prompt-based, plan-based, and learning-based methods wang2025multimodal , each contributing uniquely to the CoT generation in multimodal contexts. Prompt-based methods utilize carefully designed prompts to guide models in generating CoT during inference. Simple prompts, such as "think step-by-step" himakunthala2023let , initiate creation of reasoning paths for multimodal tasks, while advanced approaches establish a clear reasoning workflow by specifying detailed objectives and procedures for each step chen2023see ; fei2024video ; meng2023chain . Few-shot prompts often include reasoning examples chen2023grounding ; zhao2023antgpt , offering flexibility in resource-limited scenarios or when rapid responses are needed. Plan-based methods enable dynamic exploration of reasoning paths, enhancing adaptability. For example, MM-ToT kye2023mmtot combines GPT-4 hurst2024gpt and Stable Diffusion rombach2021highresolution , using depth-first and breadth-first searches to select optimal solutions. HoT yao2023thinking encapsulates interconnected thoughts within hyperedge structures, while AGoT yang2024soft builds reasoning graphs integrating visual data. Additionally, learning-based methods embed CoT construction within model training or fine-tuning. PCoT wang2024t optimizes this approach for CoT generation, whereas MC-CoT tan2024boosting improves reasoning capabilities in smaller-scale models via majority voting during training. However, as constructing medical CoT requires specific logic and domain expertise, the aforementioned methods cannot be effectively transferred to the medical field. In other words, the construction of multimodal CoT in the medical domain remains an underexplored stage. Existing studies wei2024mc primarily rely on manual annotations or model-generated reasoning paths without any evaluation. Our proposed method takes a step further by addressing these challenges by leveraging an automatic and collaborative search strategy to identify high-quality reasoning paths.
@@ 3.2 Mentor-Intern Collaborative Search for Effective Reasoning
where $\text{MICS}_{k}$ denotes the search process up to the step $k$, and $\{s_{1,x},s_{2,y},...,s_{k,z}\}$ represents the all reasoning steps searched thus far, serving as the reasoning path prefix for the subsequent search. $P$, $Q$, AA, and $I$ represent patient information, questions, ground truth, and corresponding images.
@@
$\text{MICS-Score}(s_{\leq k+1})=\frac{\text{ number of LLM($\tilde{a}_{w}$, $A$)}}{\text{ number of $\tilde{a}_{w}$}}$ (4)
@@ 4.1 Experiment Settings
In our experiments, the MICS strategy employs three mentor models: ChatGPT-4o hurst2024gpt , Gemini 2.5 Pro Preview team2024gemini , and Qwen2.5-VL-72B-Instruct bai2025qwen2 . To validate the searched reasoning paths, we designate three open-source models as intern models, each queried under two temperatures (0.3 and 1.2): Qwen2-VL-7B wang2024qwen2 , Qwen2.5-VL-7B bai2025qwen2 , and InternVL3-8B chen2024internvl . We employ DeepSeek-V3 liu2024deepseek as the “judge” to compare intern answers against the ground truth. To balance search efficiency and reliability, the maximum search depth is set to 4.
For model training, we adopt InternVL3-8B as the base model and apply LoRA fine-tuning hu2022lora with AdamW as the optimizer, a learning rate of 4e-5 with cosine decay, a batch size of 16, and bfloat16 mixed precision. Training is conducted on eight NVIDIA A100 GPUs with DeepSpeed ZeRO-1 rasley2020deepspeed . The three curriculum stages require approximately 12 hours, 12 hours, and 48 hours, respectively, totaling 72 GPU-hours. The training corpus consists of the MMRP dataset and several commonly used medical VQA datasets, with detailed composition provided in Appendix C.
@@
We compare Chiron-o1 with the following three types of baseline models: (1) General MLLMs, including high-performance general vision models, both closed-source (GPT series hurst2024gpt , Gemini series team2024gemini ) and open-source (LLaVA series liu2024improved , Qwen series bai2025qwen2 ). (2) Medical MLLMs, comprising vision models pretrained on specific medical corpora, such as LLaVA-Med li2023llava , GMAI-VL li2024gmai , and HuatuoGPT-Vision chen2024huatuogpt . (3) Medical Reasoning MLLMs, including Med-R1 lai2025med , MedVLM-R1 pan2025medvlm , and ChestX-Reasoner fan2025chestx , which are fine-tuned using GRPO shao2024deepseekmath . For evaluation metrics, we use choice accuracy to measure model performance on VQA benchmarks. For the open-ended questions requiring reasoning in Part 3 of the MMRP dataset, BERT-Score zhang2019bertscore is employed to assess the semantic similarity between the model’s final answer and the ground truth, while MICS-Score is used to evaluate the effectiveness of the reasoning path.
@@
Gemini-1.5-Pro team2023gemini 60.3 72.6 70.3 52.3 47.9 60.7
Gemini-2.5-Pro comanici2025gemini 71.3 80.5 73.9 61.1 57.1 68.8
@@
4.2 Main Results
Performance on the Medical VQA Benchmarks
Medical models should strive to further enhance their performance on VQA tasks while improving reasoning capabilities. Therefore, to evaluate the foundational performance of Chiron-o1, we first test it on several commonly-used VQA benchmarks. The results, compared against other mainstream models, are summarized in Table 1. Compared to the baseline models, Chiron-o1 achieves significant performance improvements across five benchmarks. Specifically, Chiron-o1-8B and Chiron-o1-2B outperform their respective baseline models by an average of 5.7% and 8.1%. This demonstrates that our reasoning-focused model further enhances visual understanding and question-answering capabilities. Next, we compared Chiron-o1 with SOTA models that were neither pretrained nor fine-tuned on multimodal medical datasets. The results in Table 1 indicate that our model easily surpasses these large-scale MLLMs. Subsequently, Chiron-o1 outperforms most medical MLLMs across all benchmarks. It even achieves comparable performance to HuatuoGPT-Vision-34B, which has four times the parameters, on the PMC-VQA and MMMU (H&M) benchmarks, while significantly surpassing it on the remaining benchmarks, with an overall average improvement of 5%. Finally, we observe that existing medical reasoning models, such as Med-R1 and MedVLM-R1, tend to lose basic VQA capabilities while focusing on reasoning, as shown in Table 1. In contrast, Chiron-o1-8B outperforms them on VQA benchmarks by an average of 20.6% and 16.6%, respectively. Overall, Chiron-o1 demonstrates competitive performance across benchmarks, highlighting its versatility in medical image understanding and question answering.
@@ 4.3 Ablation Studies
We apply the MICS strategy to build the MMRP multi-task dataset to improve the model’s multimodal reasoning. An ablation study on Chiron-o1-8B evaluates its impact. Unlike MICS, the vanilla method enables the mentor model to generate reasoning directly, bypassing the evaluation of intern models. Table 3 shows that excluding reasoning data (first row) or using reasoning paths constructed by vanilla method(second and third rows) degrades performance. Compared to Chiron-o1-8B, the former reduces average performance by 3% across four benchmarks, and the latter by 2.7%. Furthermore, Figure 4 illustrates the differences in effectiveness between reasoning paths constructed using MICS and the vanilla method across three distinct medical scenarios. We compute the MICS-Score step-by-step for an entire reasoning path to evaluate the value of reasoning prefixes and analyze the trend of MICS-Score changes (categorized into four types, with a monotonically increasing score indicating an effective trend and the others deemed low-value, details provided in the Appendix B). As depicted in Figure 4, the proportion of effective reasoning paths identified by MICS significantly surpasses that of the vanilla method. These results underscore the critical role of the MICS strategy in searching for effective reasoning paths.
@@
Figure 4: Ablation Studies on MICS. Contribution of the MICS strategy to reasoning path score trends, with a, b, and c denoting three clinical scenarios (Appendix A.3). "vanilla" refers to directly generating reasoning paths using the mentor model without evaluation.
Figure 5: Ablation studies on the Model Training Strategy. Figure (a) and (b) present results for Chiron-o1-8B and Chiron-o1-2B, respectively. The comparison highlights the advantage of the proposed stage-wise curriculum over alternative training schemes.
@@
During training, we incorporated VQA datasets into the training set to bolster the model’s foundational visual understanding capabilities. To evaluate their contribution to model performance, we compared a model trained without VQA datasets to Chiron-o1-8B. The results in Table 3 (second and fourth rows) indicate that VQA data significantly enhances the model’s performance on simpler benchmarks. Specifically, performance on VQA-RAD and SLAKE increased by an average of 5% and 7.3%, respectively. However, this operation has limited impact on complex and reasoning benchmarks. We further investigated whether quality control of the MMRP reasoning data could improve the model’s reasoning performance. As shown in Table 3 (fifth row), quality control enhances the model’s accuracy on MMMU (Health & Medicine) and MedXpertQA_MM, improving from 49.3% to 54.6% (+5.3%) and from 22.6% to 24.2% (+1.6%), respectively. These results effectively validate that our specific setting of the training set is both effective and reasonable.
@@
In Section 3.3, we proposed a stage-wise training strategy for fine-tuning the model. To evaluate its impact, we trained models using only subsets of the stages and compared their performance on several benchmarks against Chiron-o1-8B. First, for the complex VQA benchmark (MMMU Health & Medicine), Figure 5(a) shows that models trained with Stage 1, Stage 1+2, or Stage 3 alone exhibit significant performance gaps compared to Chiron-o1-8B, with reductions of 6.9%, 6.2%, and 3.5%, respectively. Next, we observed that omitting the third stage leads to a more pronounced degradation in reasoning capabilities. Specifically, models trained with Stage 1 or Stage 1+2 perform, on average, 3.1% and 0.25% lower than those trained with Stage 3 alone on MMMU (Health & Medicine) and MedXpertQA_MM. Similarly, training solely with Stage 3 (without reinforcing visual question-answering capabilities) severely impairs performance on VQA-RAD and SLAKE compared to other ablation models.